OpenStudy (anonymous):

Solve the equation log x = 2 log (mx) for x in terms of m.

zepdrix (zepdrix):

Hey there :) You can apply some rules of logs to get this one moving along. We can deal with the 2 by using this log rule:\[\large\rm \color{\orangered}{b\cdot \log(a)=\log(a^b)}\]

zepdrix (zepdrix):

Make sure you apply the 2 to BOTH the m and the x.

zepdrix (zepdrix):

Then you'll have something of this form: \(\large\rm \log(\text{stuff})=\log(\text{other stuff})\) And since the log function is one-to-one, we can determine that the insides must be equal!\[\large\rm stuff=other~stuff\]

OpenStudy (badhi):

another way to do this is, expand the log on the right hand side \(\log(mx) = \log(m) + \log(x)\) then solve for \(\log(x)\) and find x later

zepdrix (zepdrix):

Ya, lots of fun ways to approach logs :)

OpenStudy (irishboy123):

and none dumber than this :-) \[\log x = 2 \log (mx)\] \[\frac{\log_x x}{log_x b} = 2 \frac{\log_x mx}{log_x b}\] where b is what you started in \[1 = 2 \log_x mx\] \[x^{\frac{1}{2}} = mx \]